OpenStudy (anonymous):

For the function y = 1^x , make a table with integer values of x from 0 to 4. Then graph the function (you do not need to submit the graph). Does the graph of show exponential growth, exponential decay, or neither? Explain your thinking.

OpenStudy (anonymous):

any ideas ^^

OpenStudy (anonymous):

neither growth nor decay?

OpenStudy (anonymous):

hint: \(1^x=1\) for any \(x\)

OpenStudy (anonymous):

yeah it is a constant

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

The graph will be a straight line. 1^x will always just be 1 so the y value will always be one. This is just a graph of y=1.

OpenStudy (anonymous):

ok thanks i got it

OpenStudy (anonymous):

1^3 = 1, 1^4=1 it just straight line

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

understand?